OpenStudy (quickthing):

Linda buys blocks of chocolate that are 15 inches by 10 inches by 5 inches. She then melts the chocolate and pours it into cone molds that are 6 inches tall with a diameter of 3 inches. How many chocolate cones does she make out of each block? How much chocolate is left over?

OpenStudy (quickthing):

@amistre64

OpenStudy (confluxepic):

First you have to find the volume of the chocolate and then the volume of a cone. @quickthing

OpenStudy (quickthing):

Okay, I shall try.

OpenStudy (quickthing):

The chocolate is a cube, right?

OpenStudy (confluxepic):

The chocolate is a rectangular prism.

OpenStudy (quickthing):

Oh okay.

OpenStudy (quickthing):

let me get a picture of one.

OpenStudy (confluxepic):

A cube would have all the same sides. But this one doesn't. It is 15 inches by 10 inches by 5 inches.

OpenStudy (quickthing):

Right hang on, let me label the one I have.

OpenStudy (quickthing):

Does that look OK?

OpenStudy (confluxepic):

Yes.

OpenStudy (confluxepic):

\(\huge\ Volume=length \times base \times height\)

OpenStudy (quickthing):

Awesome. Now then, the volume of a rectangular prism is l w h

OpenStudy (quickthing):

Oh you beat me lol

OpenStudy (confluxepic):

So you just plug in the values.

OpenStudy (quickthing):

OK. 15*10*5 = 750 inches cubed

OpenStudy (confluxepic):

Correct.

OpenStudy (quickthing):

I have a few more questions. Would you take a look?

OpenStudy (quickthing):

O wait, there's another part to this.

OpenStudy (confluxepic):

Yes.

OpenStudy (quickthing):

So the molds are circles....

OpenStudy (confluxepic):

No. They are cones.

OpenStudy (quickthing):

Ohhhh.

OpenStudy (confluxepic):



OpenStudy (confluxepic):

That's how they look.

OpenStudy (quickthing):

Yes, that's right.

OpenStudy (quickthing):